What is the molar solubility (in mol L-1) of a salt with general molecular formula MX (where M is a cation and X is an anion) in a solution already containing 0.317 mol L-1 X-?

The Ksp of MX = 8.48 x 10-5 ?
Answer #1

[X-] = 0.317 M

At equilibrium:

Ksp = [M+][X-]

8.48*10^-5=(s)*(0.317+ s)

Since Ksp is small, s can be ignored as compared to 0.317

Above expression thus becomes:

8.48*10^-5=(s)*(0.317)

8.48*10^-5= (s) * 0.317

s = 2.675*10^-4 mol/L

Answer: 2.67*10^-4 mol/L